A hollow metal sphere of outside diameter of 10 cm weighing 5gm/cm3 in air floats half submerged in fresh water. Determine the radius of the cavity


1
Expert's answer
2021-11-10T10:09:13-0500

Explanations & Calculations


"\qquad\qquad
\begin{aligned}
\small w&=\small mg\\
\small &=\small \frac{4}{3}\pi(r_o^3-r_i^3)dg\\
\\
\small u&=\small w\\
\\
\small u&=\small v\rho g\\
&=\small \frac{1}{2}.\frac{4}{3}\pi r_o^3\rho g\\
\\
\small \frac{4}{3}\pi([10cm]^3-r_i^3)\times5gcm^{-3}\times g&=\small \frac{2}{3}\pi\times[10cm]^3\times 1gcm^{-3}\times g\\
\small r_i&=\small 9.7\,cm
\end{aligned}"